a typical diet has an energy content of 2400 kilocalories.if this energy is used to lift a mass of 100 kg,to what height can it be raised(Take g = 10m/s^2)

Grade:10

1 Answers

Susmita
5 years ago
Potential energy is mgh=100×10×h=1000h joule.
Given this energy is equal to 2400 KCal.=2400×1000cal=2400×1000×4.2 joule
1000h=2400×1000×4.2
Or,h=2400×4.2=10080m
Hope this helps